Damerly Company (a Utah employer) wants to give a holiday bonus check of $350 to each employee. Since it wants the check amount to be $350, it will need to gross-up the amount of the bonus. Calculate the withholding taxes and the gross amount of the bonus to be made to John Rolen if his cumulative earnings for the year are $46,910. Besides being subject to social security taxes and federal income tax (supplemental rate), a 4.95% Utah income tax must be withheld on supplemental payments.

Enter deductions beginning with a minus sign (-). You may need to adjust by a penny (subtract or add) to the gross amount to get net pay to calculate exactly to $350.00. Round your calculations and final answers to the nearest cent.

Gross bonus amount $
Federal income tax withheld
OASDI tax withheld
HI tax withheld
Utah income tax withheld
Take-home bonus check $
Answer #1

Answer-

Social Security rate 6.20%
Medicare rate 1.45%
Federal Income tax 25.00%
State income tax 4.95%
Total tax percentage 37.60%
Required gross bonus 560.90 =350 /(1-37.60%)
Gross bonus amount 560.90
Federal Income tax withheld -140.23 =560.90*25%
OASDI tax withheld -34.78 =560.90*6.2%
HI tax withheld -8.13 =560.90*1.45%
Utah Income tax withheld -27.76 =560.90*4.95%
Take home bonus check 350.00